Katie Homes and Garden Co. has 14,800,000 shares outstanding. The stock is currently selling at $82 per share. If an unfriendly outside group acquired 30 percent of the shares, existing stockholders will be able to buy new shares at 35 percent below the currently existing stock price. a. How many shares must the unfriendly outside group acquire for the poison pill to go into effect? (Do not round intermediate calculations.) Number of shares b. What will be the new purchase price for the existing stockholders? (Do not round intermediate calculations. Round your answer to 2 decimal places.) New purchase priceKatie
Answer #1
a
Number of shares = 14800000*30%= 4440000
b
New purchase price = 82*(1-0.35)= $53.30